OpenStudy (anonymous):

Regis deposits $5000 into an account for his 10-year-old child. The annual interest rate is 8.5%, compounded monthly. What monthly deposit is need to ensure the account is worth $1 million by the time the child is 55-years-old. Please explain how to get the answer.

OpenStudy (anonymous):

P=5000 D= monthly payment 1 st Month -> P+D 2 nd Month -> (P+D)(1+R)+D = P(1+R)+ D((1+R)+1) 3 rd Month --> (P+D)(1+R)+D ) (1+R) = P(1+R)^2 + D((1+R)^2 + (1+R)+1) 4 th Month --> P(1+R)^3 +D((1+R)^3+(1+R)^2 + (1+R)+1) nth Month ---> P(1+R)^(n-1) +D(1+(1+R)+..........+(1+R)^n-1)

OpenStudy (anonymous):

Using geometry sum formula \[\left(1+(1+R)+\text{...}\text{...}(1+R)^{n-1}\right)=\text{ }\frac{\left(1-(1+R)^n\right)}{1-(1+R)}\]

OpenStudy (anonymous):

substitute back \[P(1+R)^{n-1}+D\left(\frac{1-(1+R)^n}{-R}\right)\]

OpenStudy (anonymous):

Plug in number and Solve P=5000 R=.085/12 n=45*12 set it equal to 1 million and solve for D

OpenStudy (anonymous):

\[5000\left(1+\frac{.085}{12}\right)^{539}+D\left(\frac{1-\left(1+\frac{.085}{12}\right)^{540}}{-\frac{.085}{12}}\right)\text{==}1000000\] I got $124.23

OpenStudy (anonymous):

this is how it grows monthly in excel